OpenStudy (anonymous):

Really need help with this. Suppose that a population parameter is 0.1, and many samples are taken from the population. if the size of each sample is 90. which of the following will most likely be the standard deviation of the sample proportions? A. 0.095 B. 0.032 C. 0.072 D. 0.055

OpenStudy (anonymous):

what's the srd dev of a sample proportion?

OpenStudy (anonymous):

std, not srd

OpenStudy (anonymous):

\[\sigma=\sqrt{p1-p}/n\]

OpenStudy (anonymous):

plug in what you know and compute. btw it's \(\sigma =\sqrt{p\cdot q/n}\) /n inside the sqrt

OpenStudy (anonymous):

what would you expect p to be near?

OpenStudy (anonymous):

could you plug the numbers into that equation for me please?

OpenStudy (anonymous):

why can't you?

OpenStudy (anonymous):

what number would go in for q?

OpenStudy (anonymous):

q = (1-p)

OpenStudy (anonymous):

so it would be \[\sigma=\sqrt{\frac{ .1*.0 }{ 90 }}\]

OpenStudy (anonymous):

the .0 is suppose to be .9

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so the answer would be B, but really 0.031622777 summed up. Thanks man!

OpenStudy (anonymous):

you did all the work!